A ball is thrown upward from the ground with an initial speed of 16.6 m/s; at the same instant, another ball is dropped from a building 20 m high. After how long will the balls be at the same height above the ground?
Answer #1

Apply kinematic equation

the distance travelled first ball is

y1 -yo = u t - 1/2 at^2 ( yo = 0)

the distance travelled by the second ball is

y2-20 m = ut + 1/2 at^2 ( u= 0)

at y1= y2

u1t -1/2 at^2 = 20 m + 1/2 gt^2

16.6 m/s t - 1/2 * 9.8) t^2 = 20 m -1/2 (9.8) t^2

16.6 t = 20

t = 1.2 s